### v2

@@ Appendix B Auxiliary Results
Lemma 5 (Corollary 1 in Chernozhukov, Chetverikov, and Kato (2015)).
Let $(X_{1},\ldots,X_{p})^{T}$ be a centered Gaussian random vector in $\mathbb{R}^{p}$ with $\sigma_{j}^{2}\coloneq\operatorname*{\mathbb{E}}[X_{j}^{2}]>0$ for all $1\leq j\leq p$. Let $\underline{\sigma}\coloneqq\min_{1\leq j\leq p}\sigma_{j}$ and $\bar{\sigma}\coloneqq\max_{1\leq j\leq p}\sigma_{j}$. Then for every $\epsilon>0$,
@@
Lemma 6 (Lemma 8 in Chernozhukov, Chetverikov, and Kato (2015)).
@@
[…]
@@
[…]
@@
${\mathrm{P}}\left(\max_{1\leq i\leq p}X_{i}>\operatorname*{\mathbb{E}}\left[\max_{1\leq i\leq p}X_{i}\right]+\lambda\right)\leq e^{-\lambda^{2}/(2\sigma^{2})}\leavevmode\nobreak\ .$
@@
${\mathrm{P}}\left(\max_{1\leq i\leq p}X_{i}>\operatorname*{\mathbb{E}}\left[\max_{1\leq i\leq p}X_{i}\right]+\lambda\right)$ $={\mathrm{P}}\left(f(Z)-\operatorname*{\mathbb{E}}[f(Z)]>\frac{\lambda}{\sigma}\right)\leq e^{-\lambda^{2}/(2\sigma^{2})}\leavevmode\nobreak\ ,$
@@ Appendix C Additional Simulation Results
C.1 Additional Results for Simulation II
In this section, we extend the results of Simulation II from the main text by including samples of size $n=200$ and $n=1000$ together with intermediate values of $p$. A third model is also considered. It is a version of Model 2 where all individual hypotheses are violated when $\rho\neq 0$.
Model 3:
Let $X\sim Unif[-1,1]$ and $\varepsilon\sim\mathcal{N}(0,\Sigma_{\tau})$, where $\Sigma_{\tau}$ has diagonal elements equal to one and off-diagonal elements equal to $\tau$, and $\varepsilon$ is independent of $X$. Further, let
$Y_{j}=2\rho\cos(8\pi X)+\varepsilon_{j}\quad\text{ for }1\leq j\leq p.$
The results are presented in Figures 712. As expected, the rejection rates (weakly) increase with the sample size for any given combination of $p$ and $\rho$, but the qualitative patterns in Models 1 and 2 remain similar to those discussed in the main text. In Model 3, when there is no dependence between the components of $Y$, the power of the distance correlation test of Székely and Rizzo (2013) substantially improves relative to Model 2, achieving similar rejection rates to our test when $n=1000$ and $p=1000$. Its performance, however, deteriorates dramatically in the presence of dependence between the components of $Y$, while our procedure remains insensitive to such dependence in the data.
Figure 7: Simulation II for Model 1 with $\tau=0$.
@@ C.1 Additional Results for Simulation II
Figure 8: Simulation II for Model 1 with $\tau=0.5$.
@@
Figure 9: Simulation for Model 2 with $\tau=0$.
@@
Figure 10: Simulation II for Model 2 with $\tau=0.5$.
@@
Figure 11: Simulation II for Model 3 with $\tau=0$.
Figure 12: Simulation II for Model 3 with $\tau=0.5$.
Notes: The tests have a nominal level of 5%. Results are based on $S=5,000$ Monte Carlo draws. The BMB tests and the distance covariance test used $B=499$ bootstrap samples/replicates.
C.2 Simulation III
In this section, we compare our method to the independence tests of Zhou, Xu, Zhu, and Li (2024) based on Hoeffding’s D (ZXZL-D), Blum-Kiefer-Rosenblatt’s R (ZXZL-R) and Bergsma-Dassios-Yanagimoto’s $\tau^{*}$ (ZXZL-$\tau^{*}$), and to the test of Zhu, Zhang, Yao, and Shao (2020) based on an aggregated distance covariance measure (ZZYS-agg.dcov). [^5] The tests ZXZL-D, ZXZL-R, ZXZL-$\tau^{*}$, and ZZYS-agg.dcov were run using the implementation of Zhou, Xu, Zhu, and Li (2024) provided at https://github.com/Yeqing-TJ/Rank-based-test-in-high-dimension [Accessed on March 10, 2025]. Due to the high computational complexity of these additional tests, we limit the simulation setup to samples of size $n=200$ with the maximum of $p=200$ individual hypotheses. [^6] Estimation of the variance of the rank-based indices of Zhou, Xu, Zhu, and Li (2024) in one sample with $p=500$ and $n=500$ takes over 40 minutes using an Intel i7-1185G7 @ 3.00 GHz processor, which renders simulations for such settings infeasible. For comparison, our bootstrap test with $B=499$ takes less than a second to compute in this setting. Figure 13 shows that, in the considered designs, the tests of Zhou, Xu, Zhu, and Li (2024) and Zhu, Zhang, Yao, and Shao (2020) perform similarly to the distance-covariance-based tests considered in Simulation II. In particular, they exhibit higher power than our test in Model 1 when there is no dependence between the components of $Y$, but they performance deteriorates significantly in the presence of such dependence. They also have virtually no power against the cosine alternatives in Models 2 and 3.
Figure 13: Simulation III.
Notes: The tests have nominal level of 5%. Results for $n=200$ and $S=1,000$ Monte Carlo draws. The BMB tests and the distance covariance used $B=499$ bootstrap samples/replicates. Where not distinguishable, the power curves of the tests dcorT.test, dcov.test, ZXZL-D, ZXZL-R, ZXZL-$\tau^{*}$, and ZZYS-agg.cov overlap.

### v3

@@ Appendix B Auxiliary Results
Lemma 5 (Corollary 1 in Chernozhukov et al. (2015)).
Let $(X_{1},\ldots,X_{p})^{T}$ be a centred Gaussian random vector in $\mathbb{R}^{p}$ with $\sigma_{j}^{2}\coloneq\operatorname*{\mathbb{E}}[X_{j}^{2}]>0$ for all $1\leq j\leq p$. Let $\underline{\sigma}\coloneqq\min_{1\leq j\leq p}\sigma_{j}$ and $\bar{\sigma}\coloneqq\max_{1\leq j\leq p}\sigma_{j}$. Then for every $\epsilon>0$,
@@
Lemma 6 (Lemma 8 in Chernozhukov et al. (2015)).
@@
Let $X_{1},\dots,X_{n}$ be independent random variables in $\mathbb{R}^{k}$ with CDF $F(\cdot)$, and let $\hat{F}$ be their empirical CDF . There exist constants $c_{k}$ and $c_{k}^{\prime}$ such that for all $\varepsilon>0$ and $n\in\mathbb{N}$ it holds that
@@
Suppose that $n=2N$, and let $F_{n}^{\prime}$ be the empirical CDF calculated based on observations with odd indices $i$, and $F_{n}^{\prime\prime}$ based on the even indices. Then for any $\varepsilon>0$, using the union bound, we obtain that
@@
${\mathrm{P}}\left(\max_{1\leq i\leq p}X_{i}>\operatorname*{\mathbb{E}}\left[\max_{1\leq i\leq p}X_{i}\right]+\lambda\right)\leq e^{-\lambda^{2}/(2\sigma^{2})}\penalty 10000\ .$
@@
${\mathrm{P}}\left(\max_{1\leq i\leq p}X_{i}>\operatorname*{\mathbb{E}}\left[\max_{1\leq i\leq p}X_{i}\right]+\lambda\right)$ $={\mathrm{P}}\left(f(Z)-\operatorname*{\mathbb{E}}[f(Z)]>\frac{\lambda}{\sigma}\right)\leq e^{-\lambda^{2}/(2\sigma^{2})}\penalty 10000\ ,$
@@ Appendix C Additional Simulation Results
C.1 Additional Results for Experiment 1
Figure 10 complements the results from the main text by considering the case where $Y_{1},\ldots,Y_{p}$ have baseline correlation. The results are very similar to those presented in Figure 3, which shows that our procedure is not sensitive to dependence between the individual test statistics.
Figure 10: Rejection rates in Models 1–6 with $\tau=0.5$ under different choices of the big block size $q$.
Notes: The tests have nominal level of 5%. The null of independence holds if $\rho=0$, while $\tau$ denotes the correlation between variables $Y_{1},\ldots,Y_{p}$ under the null. Results for sample size $n=500$, $B=499$ bootstrap replications, and $S=5,000$ Monte Carlo draws.
C.2 Additional Results for Experiments 2 and 4
In this section, we extend the results of Experiments 2 and 4 from the main text by including samples of size $n=200$ and $n=1000$. The results are presented in Figures 1114. As expected, the rejection rates (weakly) increase with the sample size for any given combination of $p$ and $\rho$, but the qualitative patterns remain similar to those discussed in the main text.
Figure 11: Experiments 2 and 4 for linear alternatives with $n=200$.
@@ C.2 Additional Results for Experiments 2 and 4
Figure 12: Experiments 2 and 4 for linear alternatives with $n=1000$.
@@
Figure 13: Experiments 2 and 4 for cosine alternatives with $n=200$.
@@
Figure 14: Experiments 2 and 4 for cosine alternatives with $n=1000$.